Suppose a particular tissue culture has area AA(tt) at time tt and a

potential maximum area MM. Based on properties of cell division, it is reasonable to assume that

the area AA grows at a rate jointly proportional to �AA(tt) and MM − AA(tt); that is


dddd

dddd = kk�AA(tt) [MM − AA(tt)]


where kk is a positive constant.


1 You may wish to begin your research by consuming the following articles:

Robert Eisner, The Misunderstood Economics: What Counts and How to Count It, Boston, MA: Harvard Business School Press, 1994, pp.

196-199 and Robert H. Frank, Microeconomics and Behavior, 2nd ed., New York: McGraw-Hill, 1994, pp. 656-657.


4 | P a g e


a. Let RR(tt) = AA′


(tt) be the rate of tissue growth. Show that RR′


(tt) = 0 when AA(tt) = MM/3.

b. Is the rate of tissue growth greatest or least when AA(tt) = MM/3? [Hint: Use the first

derivative test or second derivative test.]

c. Based on the given information and what you discovered in part (a), what can you say

about the graph of AA(tt)?


1
Expert's answer
2021-11-30T11:34:36-0500
"\dfrac{dA}{dt}=k\sqrt{A(t)}(M-A(t))"

a.


"R(t)=A'(t)"

"R'(t)=A''(t)=\bigg(k\big(M\sqrt{A(t)}-A(t)\sqrt{A(t)}\big)\bigg)'"

"=k(\dfrac{M}{2\sqrt{A}}-\dfrac{3\sqrt{A}}{2})A'"

"=\dfrac{k}{2\sqrt{A}}(M-3A)\big(k\sqrt{A}(M-A)\big)"

"=\dfrac{k^2}{2}(M-3A)(M-A)"

"R'(t)=0=>\dfrac{k^2}{2}(M-3A)(M-A)=0"

"A=\dfrac{M}{3}\ or\ A=M"


b.

If "A<\dfrac{M}{3}, R'>0, R" increases.


If "\dfrac{M}{3}<A<M, R'<0, R" decreases.

The rate of tissue growth "R(t)=A'(t)" is greatest when "A(t)=\dfrac{M}{3}."


c.

"A''(t)=0," ​when"A=\dfrac{M}{3}."

If "A<\dfrac{M}{3}, A''>0, A" is concave up.


If "\dfrac{M}{3}<A<M, A''<0, A" is concave down.


The graph of "A(t)" has an inflection point when "A(t)=\dfrac{M}{3}."